OpenStudy (anonymous):

If i is the imaginary unit such that sqrt(-1) = i, then (7i)2 = ?

OpenStudy (anonymous):

(7i)^2

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

See it as 7 multiplied by i all squared If you prefer xy all squared which is x^2y^2 Therefore 7i all squared is 7^2i^2 which is 49 x -1 - -49

OpenStudy (anonymous):

-49

OpenStudy (anonymous):

-1^2=-1?

OpenStudy (anonymous):

no of course not. it is \[i^2=-1\] because \[i=\sqrt{-1}\]

OpenStudy (anonymous):

no i^2 = -1